OpenStudy (anonymous):

a force of 300 lbs slow down a wagon from 60 ft/s to 20 ft/s in 10 s. what will be the wagon's weight?

OpenStudy (ghazi):

since F=m*a now to find a you need to write V^2=U^2+a*t so (18.75)^2=(6.25)^2+a*10 , now 1m = 3.2 ft so 60ft/s=18.75m/s and 20ft/s = 6.25m/s substitute this value in above equation we get a=31.25 ....now F=300lbs = 140.7 N= m* 31.25 ...and you'll get mass= 4.5kg

OpenStudy (anonymous):

we know that,